# On $\ell$-regular partitions and Hickerson’s identity

### v1

@@
Abstract. Based on two involutions and a bijection, we completely determine the difference of the number of $\ell$-regular partitions of $n$ into an even number of parts and into an odd number of parts for all positive integers $\ell>1$, which extends two recent results due to Ballantine and Merca. As an application, we provide a combinatorial proof of Hickerson’s identity on the number of partitions into an even and odd number of parts.
@@ 1 Introduction
The first aim of the paper is to determine $b_{\ell}^{e}(n)-b_{\ell}^{o}(n)$ for all positive integers $\ell$ and $n$ with $\ell>1$ through a combinatorial approach, which extends (1.1) and (1.2).
@@
Notice that Hickerson’s proof of (1.7) relies on the generating functions. The second aim of the paper is to provide a combinatorial proof of (1.7) by using (1.4). We shall present combinatorial proofs of (1.3), (1.4) and (1.7) in Sections 2–4, respectively.
@@
For an $\ell$-regular partition $\lambda$ of $n$, let $e_{\ell}(\lambda)$ denote the largest even part of $\lambda$ which appears an odd number of times, and $t_{\ell}(\lambda)$ denote the largest repeated part of $\lambda$ which is not congruent to $\ell/2\pmod{\ell}$. For the sake of convenience, let $e_{\ell}(\lambda)=0$ for a partition $\lambda$ with all even parts occurring an even number of times, and $t_{\ell}(\lambda)=0$ for a partition $\lambda$ with no repeated part $\not\equiv\ell/2\pmod{\ell}$. Let $B^{\prime}_{\ell}(n)$ be the set of $\ell$-regular partitions $\lambda$ of $n$ with $e_{\ell}(\lambda)\not=0$ or $t_{\ell}(\lambda)\not=0$. We define the involution $\psi_{\ell}$ on $B^{\prime}_{\ell}(n)$ as follows.
@@
Let $Q_{r}(n)$ be the set of partitions of $n$ where each part occurs at most $r$ times, and $B_{\ell}(n)$ be the set of $\ell$-regular partitions of $n$. We shall establish a bijection $\varphi_{r}:Q_{r}(n)\to B_{r+1}(n)$.

### v2

@@
Abstract. Based on two involutions and a bijection, we completely determine the difference between the number of $\ell$-regular partitions of $n$ into an even number of parts and into an odd number of parts for all positive integers $n$ and $\ell>1$, which extends two recent results due to Ballantine and Merca. As an application, we provide a combinatorial proof of Hickerson’s identity on the number of partitions into an even and odd number of parts.
@@ 1 Introduction
Inspired by Ballantine and Merca’s combinatorial approach to (1.1) and (1.2), we completely determine $b_{\ell}^{e}(n)-b_{\ell}^{o}(n)$ for all positive integers $n$ and $\ell>1$ through two involutions and a bijection, which extends (1.1) and (1.2).
@@
Notice that Hickerson’s proof of (1.7) relies on the generating functions. The second aim of the paper is to provide a combinatorial proof of (1.7) by using (1.4) and a bijection due to Glaisher [6]. We shall present combinatorial proofs of (1.3), (1.4) and (1.7) in Sections 2–4, respectively.
@@
For an $\ell$-regular partition $\lambda$ of $n$, let $e_{\ell}(\lambda)$ denote the largest even part of $\lambda$ which appears an odd number of times, and $t_{\ell}(\lambda)$ denote the largest repeated part of $\lambda$ which is not congruent to $\ell/2$ modulo $\ell$. For the sake of convenience, let $e_{\ell}(\lambda)=0$ for a partition $\lambda$ with all even parts occurring an even number of times, and $t_{\ell}(\lambda)=0$ for a partition $\lambda$ with no repeated part $\not\equiv\ell/2\pmod{\ell}$. Let $B^{\prime}_{\ell}(n)$ be the set of $\ell$-regular partitions $\lambda$ of $n$ with $e_{\ell}(\lambda)\not=0$ or $t_{\ell}(\lambda)\not=0$. We define the involution $\psi_{\ell}$ on $B^{\prime}_{\ell}(n)$ as follows.
@@
Let $Q_{r}(n)$ be the set of partitions of $n$ where each part occurs at most $r$ times, and $B_{\ell}(n)$ be the set of $\ell$-regular partitions of $n$. There is a classic bijection between $Q_{r}(n)$ and $B_{r+1}(n)$ due to Glaisher [6], which is stated as follows.